OpenStudy (anonymous):

723 tickets were sold for $3 to students and $5 to non students. They raised $2815. How many tickets were sold to students? to non-students?

OpenStudy (anonymous):

@satellite73 Before you do this, please come help me really quick

OpenStudy (anonymous):

i guess you are supposed to solve \[3x+5y=2815\\x+y=723\]

OpenStudy (anonymous):

in the second equation you can put \(y=723-x\) and substitute in to the first equation to get \[3x+5(723-x)=2815\] and solve that one

OpenStudy (anonymous):

Thank you, so x will be the number of students and 723-x is th non-student tickets right?